The number of times the AI algorithm is successful at detecting fake news is normally distributed with a sample mean of 900 and the sample standard deviation of 30. Assume a sample osize of 100 was used.
What is the 99% confidence interval estimate for the population mean?
1. (862.6636 ; 937.3364)
2. (864.2155 ; 935.7845)
3. (892.1208 ; 907;8792)
4. (892.9062 ; 907;0938)
5. (892.2600 ; 907.7400)
1
Expert's answer
2020-09-24T16:47:54-0400

We need to construct the 99% confidence interval for the population mean 

μ. The following information is provided:


sample mean "\overline{X}=900\\s=30\\N=100"


since sample size is larger and normally distributed , we use Z confidence interval .


The critical value for α=0.01 is "z_c = z_{1-\alpha/2} = 2.576"

.The corresponding confidence interval is computed as shown below:

"\\begin{array}{ccl} CI= \\displaystyle \\left( \\bar X - z_c \\times \\frac{\\sigma}{\\sqrt{n}}, \\bar X + z_c \\times \\frac{\\sigma}{\\sqrt{n}} \\right) \\\\ \\\\ \\displaystyle= \\left( 900 - 2.576 \\frac{30}{\\sqrt{100}} , 900 + 2.576 \\frac{30}{\\sqrt{100}} \\right) \\\\ \\\\ = (892.26, 907.74) \\end{array}\n\n\n\n\n\n\n\n\n\u200b"

Answer: option 5